# Polynomial-Time Preparation of Low-Temperature Gibbs States for 2D Toric Code

### v1

@@ Abstract
We propose a polynomial-time algorithm for preparing the Gibbs state of the two-dimensional toric code Hamiltonian at any temperature, starting from any initial condition, significantly improving upon prior estimates that suggested exponential scaling with inverse temperature. Our approach combines the Lindblad dynamics using a local Davies generator with simple global jump operators to enable efficient transitions between logical sectors. We also prove that the Lindblad dynamics with a digitally implemented low temperature local Davies generator is able to efficiently drive the quantum state towards the ground state manifold. Despite this progress, we explain why protecting quantum information in the 2D toric code with passive dynamics remains challenging.
@@ 1 Introduction
In this work, we propose a novel Gibbs sampler with nonlocal jump operators, and analyze its convergence rate for preparing low-temperature Gibbs states of the 2D toric code [33], a paradigmatic model in quantum information theory, quantum error correction, and condensed matter physics. In the context of thermalization, when the toric code is exposed to thermal noise modeled by a specific form of Lindbladians called the Davies generator, the seminal work of Alicki et al. [2] showed that the inverse spectral gap (which leads to an upper bound of the mixing time) grows exponentially with inverse temperature $\beta$. Using a different argument based on energy barriers, Temme et al. [50, 49] confirmed that the thermalization time (i.e., mixing time of the Lindblad dynamics) of the 2D toric code with local noise should indeed scale exponentially with $\beta$. However, it is unknown whether the $\exp(\beta)$ factor in the mixing time is unavoidable for all Lindblad dynamics on this problem. This leads to the central question of this work:
Can we design a quantum algorithm, based on Lindblad dynamics, that efficiently prepares the low-temperature Gibbs state of the 2D toric code with a polynomial runtime in both the inverse temperature $\beta$ and the number of qubits $N$?
@@ 1.1 Contribution
$\mathcal{L}_{\beta}=\underbrace{\sum^{N}_{j=1}\mathcal{L}_{\sigma^{x}_{j}}+\mathcal{L}_{\sigma^{y}_{j}}+\mathcal{L}_{\sigma^{z}_{j}}}_{:=\mathcal{L}_{\text{\rm local full}}}+\underbrace{\mathcal{L}_{\mathsf{X}_{1}}+\mathcal{L}_{\mathsf{Z}_{1}}+\mathcal{L}_{\mathsf{X}_{2}}+\mathcal{L}_{\mathsf{Z}_{2}}}_{:=\mathcal{L}_{\rm global}}\,,$ (1.1)
@@
${\rm Gap}(-\mathcal{L}_{\beta})=\max\left\{e^{-\mathcal{O}(\beta)},\Omega(N^{-3})\right\}\,.$
@@
The choice of the global jump operators in (1.1) is natural. As $\beta\to\infty$, the thermal state converges toward the ground state of the Hamiltonian. Since the 2D toric code has four degenerate ground states that are not locally connected, the local Davies generator $\mathcal{L}_{\text{\rm local full}}$ cannot efficiently transit between these ground states, resulting in a slow mixing process. To overcome this difficulty, we introduce the global jump operators $\mathcal{L}_{\rm global}$ that enable transitions between different ground states at low temperatures, which ensures fast mixing even at zero temperature. This phenomenon is not unique to the 2D toric code. For example, in our paper, we also consider a simpler 1D Ising model in Appendix A, construct a similar Davies generator with global jump operators, and demonstrate a fast mixing result similar to 1 to illustrate the proof concept.
@@
$\max\left\{\exp(-\mathcal{O}(\beta)),\Omega(N^{-3})\right\}\,.$
The above proposition implies that, even in the absence of the global jump operator, the system thermalizes quickly within the syndrome space. This is a key step towards proving 1; see Section 3 for further details. Here the operator $\mathcal{L}_{\text{local}}$ in 2 is carefully designed so that analyzing syndrome mixing reduces to studying the spectral gap of two distinct Glauber dynamics with quasi-1D classial Ising Hamiltonians (precisely, on the “snake” and “comb”, see Fig. 3) at low temperature. We then introduce a new iterative method to establish lower bounds of their spectral gaps. To the best of our knowledge, this result concerning low-temperature thermalization for such quasi-1D classical Ising models is also novel in the literature.
@@ 1.2 Implications
$t_{\rm mix}(\epsilon)=\mathcal{O}\left(\frac{1}{\alpha}\left(N+\beta\right)\right)\,.$
@@
By substituting the spectral gap estimate from 1, the mixing time of $\mathcal{L}_{\beta}$ in Eq. 1.1 scales as $\mathcal{O}\left(\min\left\{\beta\mathrm{poly}(N),\exp(c\beta)(N+\beta)\right\}\right)$ for some universal constant $c$, which is a significant improvement over the $\mathcal{O}\left(\exp(c\beta)(N+\beta)\right)$ bound given in [2] when $\beta\gg 1$. It is worth mentioning that in both cases, the mixing time scales polynomially with the number of qubits, $N$, which is referred to as fast mixing in the literature. In our case, the $N$ dependence arises from both the spectral gap and the prefactor $\max_{\rho}\chi^{2}(\rho,\sigma_{\beta})$. While the latter dependence can sometimes be improved to $\mathcal{O}(\log(N))$ by considering the relative entropy distance [5, 6, 30], achieving the rapid mixing regime. This would also require translating the spectral gap lower bound into a constant or $\Omega(\log(N)^{-1})$ lower bound for the modified logarithmic Sobolev inequality (MLSI) constant. This presents an interesting direction for future exploration.
Thermal state versus ground state preparation:
@@
The 2D toric code is a stabilizer Hamiltonian, and its ground state can be efficiently prepared by measuring all stabilizers. The ground state can also be prepared using carefully designed dissipative state engineering approaches (see e.g. [53]). However, these approaches do not easily generalize to algorithms for thermal state preparation. The fast mixing result at low temperatures in this work suggests that thermal state preparation methods can also serve as a generic tool for approximate ground state preparation by selecting a sufficiently low temperature. Specifically, as discussed in 2, the local Davies generator within the syndrome space has a spectral gap that decays only polynomially with system size as $\beta$ approaches infinity. This implies that at fixed sufficiently low temperature ($\beta\gg 1$), once a quasi-particle pair (an elementary excited state of the 2D toric code, see Section 4.1) appears in the system, the local Davies generator can eliminate it in $\mathcal{O}(\text{poly}(N))$ time. Consequently, if the goal is to efficiently prepare some ground state while discarding logical information, it is sufficient to use local Davies generators, and the time required is much shorter than that for equilibrating all logical sectors using local Davies generators. This result may be of independent interest, as nature itself often prepares ground states by cooling.
Fast thermalization from a ground state:
@@
A natural question regarding the thermalization of the toric code Hamiltonian is: if one starts from the ground state, doesn’t it take exponential time in $\beta$ to create even one quasi-particle pair in the first place? While this is correct, and may seem paradoxical, it does not contradict our main result that the thermalization time scales polynomially in $\beta$. The reason is that although creating one quasi-particle pair from the ground state takes an exponential amount of time in $\beta$ to create one quasi-particle pair from the ground state, the fraction of the excited state in the thermal state is exponentially small in $\beta$. In the case of 2D toric code, the slow thermalization of local Davies generators at low temperatures mainly stems from the need to achieve an equal population across all ground states. In other words, the challenge of thermalization lies mainly in transitioning between orthogonal ground states, as discussed in Section 1.1. In our work, we prove that spectral gap of the global jump operators restricted to the logical space is, independent of $\beta$ at low temperatures. This allows the system to equilibrate rapidly, even starting from a ground state.
@@
Fast thermal state preparation versus quantum memory:
@@
A good self correcting quantum memory (SCQM) should be able to store a quantum state in contact with a cool thermal bath for a duration that increases exponentially with the size of the system. If the thermalization time of a quantum system only scales polynomially with system size, it cannot be considered a viable candidate for SCQM. Ref. [2] demonstrated that, at any constant temperature, the 2D toric code has a spectral gap that remains independent of system size, and thus is not a good SCQM, and to date valid candidates for SCQM are only known in 4D or higher [3]. Our refined estimate indicates that local Davies generator at low temperature (or even at zero temperature) can be efficient in annihilating quasi-particles in the syndrome space.
@@
A natural question is: does it make the 2D toric code a candidate for passively protected quantum memories (sometimes known as autonomous quantum error correction, autonomous quantum memory protection) [13, 37, 38]? Specifically, consider a Lindbladian $\mathcal{L}=\mathcal{L}_{e}+\mathcal{L}_{r}$ where $\mathcal{L}_{e}$ is a Davies generator modeling thermal noise at some finite temperature $\beta^{-1}$, and $\mathcal{L}_{r}$ is a Davies generator at near-zero temperature (implemented either digitally or using analogue devices), and the number of terms (each of up to unit strength) in $\mathcal{L}_{r}$ can scale polynomially in $N$. Starting from a pure ground state $\rho_{0}$ carrying well-defined logical information, we run the dynamics for some fixed time $t$, and then apply a single round of decoding map $\mathcal{E}_{d}$ to obtain a final density matrix $\rho_{f}$. Can we ensure that the trace distance between $\rho_{0}$ and $\rho_{f}$ decreases super-polynomially in $N$? Unfortunately, the answer is very likely negative for the 2D toric code. The reason is that once $\mathcal{L}_{e}$ creates a syndrome in the form of a quasi-particle excitation, this quasi-particle can be diffused by either $\mathcal{L}_{e}$ or $\mathcal{L}_{r}$ at zero energy cost. As a result, in the worst case, a quasi-particle may diffuse across the torus in polynomial time before it is annihilated, which changes the logical information. In this sense, the key difference between protecting logical information through passive dynamics and using quantum error correction is that the latter actively guides the diffusion and annihilation of quasi-particles in a way that preserves the logical information.
Notably, this problem does not arise in 2D Ising models for which local excitations cannot diffuse without incurring an energy cost. This is a key aspect in recent designs of passively protected quantum memories [38]. To our knowledge, the theoretical analysis of such models remains an open question.
@@ 1.3 Related works
The thermalization of stabilizer Hamiltonians using a local Davies generator has been explored in several prior works [2, 50, 49, 22, 11, 23]. In [2], the authors demonstrated that the local Davies generator achieves fast thermalization for the 2D toric code. In particular, the spectral gap of the Davies generator, when considering all local Pauli coupling operators, is lower bounded by $\exp(-\Theta(\beta))$. The bound is valid for all temperatures and results in a mixing time (defined via the trace distance) scaling as $t_{\rm mix}=N\exp(\Theta(\beta))$. Along the same direction, [50, 49] considered general stabilizer codes and establish a lower bound on the spectral gap (or Poincaré constant) using the generalized energy barrier $\overline{\epsilon}$ [49, Definition 13]. Specifically, for any given $\beta>0$, the spectral gap can be lower bounded by $C_{N}\exp\left(-\beta\overline{\epsilon}\right)$, where $C_{N}$ is a technical constant that typically scales as $1/N$. Although these works address thermalization across all temperatures, these lower bounds on the spectral gap are insufficient for efficiently preparing low-temperature thermal states, as the gap decays exponentially in $\beta$. Specifically, when using a local Davies generator to transition between ground states along a local Pauli path, the energy must first increase, requiring the dynamics to overcome the energy barrier to fully thermalize. Furthermore, as analyzed in [49, 34], an exponentially small spectral gap of order $\exp(-\Theta(\beta))$ in a local Davies generator appears inevitable when such energy barriers are present.
In our work, to overcome the bottleneck posed by the energy barrier, we modify the local Davies generator by incorporating appropriate global coupling operators that can directly connect the degenerate ground states, thereby avoiding the issues associated with the generalized energy barrier defined by local Pauli paths. By refining the analysis in [2], we demonstrate that the resulting dynamics exhibits a spectral gap that decays polynomially with the size of the system but remains independent of $\beta$, ensuring fast mixing even at low temperatures. We note that the polynomially decaying spectral gap primarily arises from the mixing rate within the syndrome space, which remains unaffected by the introduction of the new global jump operators acting on the logical space. Consequently, our refined mixing time estimate rigorously demonstrates that the local Davies generator at low temperatures (and even at zero temperature) can efficiently annihilate quasi-particles residing in the syndrome space. A similar phenomenon has been numerically observed in the study of quantum memory [22, 11, 23].

### v2

@@ Abstract
We propose a polynomial-time algorithm for preparing the Gibbs state of the two-dimensional toric code Hamiltonian at any temperature, starting from any initial condition, significantly improving upon prior estimates that suggested exponential scaling with inverse temperature. Our approach combines the Lindblad dynamics using a local Davies generator with simple global jump operators to enable efficient transitions between logical sectors. Our proof also shows that the Lindblad dynamics with a digitally implemented low-temperature local Davies generator is able to efficiently drive the quantum state towards the ground state manifold.
@@ 1 Introduction
In this work, we propose a novel Gibbs sampler with nonlocal jump operators, and analyze its convergence rate for preparing low-temperature Gibbs states of the 2D toric code [36], a paradigmatic model in quantum information theory, quantum error correction, and condensed matter physics. In the context of thermalization, when the toric code is exposed to thermal noise modeled by a specific form of Lindbladians called the Davies generator, the seminal work of Alicki, Fannes, and Horodecki [2] showed that the inverse spectral gap (which leads to an upper bound of the mixing time) grows exponentially with inverse temperature $\beta$. Using a different argument based on energy barriers, Temme and Kastoryano [53, 52] confirmed that the thermalization time (i.e., mixing time of the Lindblad dynamics) of the 2D toric code with local noise should indeed scale exponentially with $\beta$. However, it is unknown whether the $\exp(\beta)$ factor in the mixing time is unavoidable for all Lindblad dynamics on this problem. This leads to the central question of this work:
Can we design a quantum algorithm, based on Lindblad dynamics, that efficiently prepares the low-temperature Gibbs state of the 2D toric code from any initial state with a polynomial runtime in both the inverse temperature $\beta$ and the number of qubits $N$?
@@
$\sigma^{x}=\begin{bmatrix}0&1\\ 1&0\end{bmatrix},\quad\sigma^{y}=\begin{bmatrix}0&-i\\ i&0\end{bmatrix},\quad\sigma^{z}=\begin{bmatrix}1&0\\ 0&-1\end{bmatrix}.$
@@ 1.1 Contribution
$\mathcal{L}_{\beta}=\underbrace{\sum^{N}_{j=1}\left(\mathcal{L}_{\sigma^{x}_{j}}+\mathcal{L}_{\sigma^{y}_{j}}+\mathcal{L}_{\sigma^{z}_{j}}\right)}_{:=\mathcal{L}_{\text{\rm local full}}}+\underbrace{\mathcal{L}_{\mathsf{X}_{1}}+\mathcal{L}_{\mathsf{Z}_{1}}+\mathcal{L}_{\mathsf{X}_{2}}+\mathcal{L}_{\mathsf{Z}_{2}}}_{:=\mathcal{L}_{\rm global}}\,,$ (1.1)
@@
${\rm Gap}(-\mathcal{L}_{\beta})=\max\left\{\exp(-\mathcal{O}(\beta)),\Omega(N^{-2})\right\}\,.$
@@
The choice of the global jump operators in (1.1) is natural. As $\beta\to\infty$, the thermal state converges toward the ground state of the Hamiltonian. Since the 2D toric code has four degenerate ground states that are not locally connected, the local Davies generator $\mathcal{L}_{\text{\rm local full}}$ cannot efficiently transit between these ground states, resulting in a slow mixing process. To overcome this difficulty, we introduce the global jump operators $\mathcal{L}_{\rm global}$ that enable transitions between different ground states at low temperatures, which ensures fast mixing even at zero temperature. This phenomenon is not unique to the 2D toric code. For example, in our paper, we also consider a simpler 1D Ising model in Appendix B, construct a similar Davies generator with global jump operators, and demonstrate a fast mixing result similar to 1 to illustrate the proof concept.
@@
$\max\left\{\exp(-\mathcal{O}(\beta)),\Omega(N^{-2})\right\}\,.$
The above proposition implies that, even in the absence of the global jump operator, the system thermalizes quickly within the syndrome space. This is a key step towards proving 1; see Section 2 for further details. Here the operator $\mathcal{L}_{\text{local}}$ in 2 is carefully designed so that analyzing syndrome mixing reduces to studying the spectral gap of two distinct Glauber dynamics with quasi-1D classical Ising Hamiltonians (precisely, on the “snake” and “comb”, see Fig. 3) at low temperature. Then, to derive lower bounds of their spectral gaps, we introduce a new iterative method and establish a new tight estimate on the minimal eigenvalue of the perturbed graph Laplacian of the stair graph. To the best of our knowledge, this result concerning low-temperature thermalization for such quasi-1D classical Ising models is also novel in the literature.
@@ 1.2 Implications
$t_{\rm mix}(\epsilon)=\mathcal{O}\left({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{N}{\alpha}\left(1+\beta\right)}\right)\,.$
@@
By substituting the spectral gap estimate from 1, the mixing time of $\mathcal{L}_{\beta}$ in Eq. 1.1 scales as $\mathcal{O}\left(\min\left\{\beta\mathrm{poly}(N),\exp(c\beta)(N+\beta)\right\}\right)$ for some universal constant $c$, which is a significant improvement over the $\mathcal{O}\left(\exp(c\beta)(N+\beta)\right)$ bound given in [2] when $\beta\gg 1$. It is worth mentioning that in both cases, the mixing time scales polynomially with the number of qubits $N$, which is referred to as fast mixing in the literature. In our case, the $N$ dependence arises from both the spectral gap and the prefactor $\max_{\rho}\chi^{2}(\rho,\sigma_{\beta})$. While the latter dependence can sometimes be improved to $\mathcal{O}(\log(N))$ by considering the relative entropy distance [5, 6, 33], achieving the rapid mixing regime. This would also require translating the spectral gap lower bound into a constant or $\Omega(\log(N)^{-1})$ lower bound for the modified logarithmic Sobolev inequality (MLSI) constant. This presents an interesting direction for future exploration.
Thermal state versus ground state preparation.
The 2D toric code is a stabilizer Hamiltonian, and its ground state can be efficiently prepared using local measurements. Recently, Hwang and Jiang [30] developed an efficient quantum protocol for preparing Gibbs states of the 2D toric code in $\mathcal{O}(N^{2})$ time uniformly in any positive temperature, and the approach can be generalized to the defected toric code. It constructs quantum Gibbs states by reducing the problem to Gibbs sampling of a classical local Hamiltonian. The goal of these protocols is to prepare ground or thermal states, starting from specific states such as tensor product states or maximally mixed states. In contrast, the goal of this work is different: it investigates whether dissipative dynamics can drive any initial state toward thermal equilibrium, analogous to certain dissipative state engineering approaches used for ground state preparation (see, e.g., [56]).
@@
However, these approaches do not easily generalize to algorithms for thermal state preparation. The fast mixing result at low temperatures in this work suggests that thermal state preparation methods can also serve as a generic tool for ground state preparation by selecting a sufficiently low temperature. Specifically, as discussed in 2, the local Davies generator within the syndrome space has a spectral gap that decays only polynomially with system size as $\beta$ approaches infinity. This implies that at fixed sufficiently low temperature ($\beta\gg 1$), once a quasi-particle pair (an elementary excited state of the 2D toric code, see Section 4.1) appears in the system, the local Davies generator can eliminate it in $\text{poly}(N)$ time. Consequently, if the goal is to efficiently prepare some ground state while discarding logical information, it is sufficient to use local Davies generators, and the time required is much shorter than that for equilibrating all logical sectors using local Davies generators.
Fast thermalization from a ground state.
@@
A natural question regarding the thermalization of the toric code Hamiltonian is: if one starts from the ground state, doesn’t it take exponential time in $\beta$ to create even one quasi-particle pair in the first place? While this is correct and may seem paradoxical, it does not contradict our main result that the thermalization time scales polynomially in $\beta$. The reason is that although creating one quasi-particle pair from the ground state takes an exponential amount of time in $\beta$, the fraction of the excited state in the thermal state is exponentially small in $\beta$. In the case of 2D toric code, the slow thermalization of local Davies generators at low temperatures mainly stems from the need to achieve an equal population across all ground states. In other words, the challenge of thermalization lies mainly in transitioning between orthogonal ground states, as discussed in Section 1.1. In our work, we prove that the spectral gap of the global jump operators restricted to the logical space is independent of $\beta$ at low temperatures. This allows the system to equilibrate rapidly, even starting from a ground state.
@@
Fast thermal state preparation, quantum memory, and cellular automaton.
@@
A good self-correcting quantum memory (SCQM) should be able to store a quantum state in contact with a cool thermal bath for a duration that increases exponentially with the size of the system. This is sometimes also referred to as passively protected quantum memories, autonomous quantum error correction, or autonomous quantum memory protection [14, 40, 41]. If the thermalization time of a quantum system only scales polynomially with system size, it cannot be considered a viable candidate for SCQM. [2] demonstrated that at any constant temperature, the 2D toric code has a spectral gap that remains independent of system size and thus is not a good SCQM, and to date, valid candidates for SCQM are only known in 4D or higher [3]. Our refined estimate indicates that local Davies generators at low temperatures (or even at zero temperature) can be efficient in annihilating quasi-particles in the syndrome space.
@@
However, can we digitally implement a low-temperature Davies generator to protect the logical information in the 2D toric code? Specifically, consider a Lindbladian $\mathcal{L}=\mathcal{L}_{e}+\mathcal{L}_{r}$ where $\mathcal{L}_{e}$ is a Davies generator modeling thermal noise in nature at some finite temperature $\beta^{-1}$, and $\mathcal{L}_{r}$ is a digitally implemented Davies generator at near-zero temperature, and the number of terms (each of up to unit strength) in $\mathcal{L}_{r}$ can scale polynomially in $N$. Starting from a pure ground state $\rho_{0}$ carrying well-defined logical information, we run the dynamics for some fixed time $t$, and then apply a single round of decoding map $\mathcal{E}_{d}$ to obtain a final density matrix $\rho_{f}$. Can we ensure that the trace distance between $\rho_{0}$ and $\rho_{f}$ decreases super-polynomially in $N$? This procedure is similar to continuous-time recovery with time-independent control (also called continuous-time quantum error correction; see e.g., [19, 38]). This task appears to be highly challenging. The reason is that once $\mathcal{L}_{e}$ creates a syndrome in the form of a quasi-particle excitation, this quasi-particle can be diffused by either $\mathcal{L}_{e}$ or $\mathcal{L}_{r}$ at zero energy cost. As a result, in the worst case, a quasi-particle may diffuse across the torus in polynomial time before it is annihilated, which changes the logical information.
The problem of dissipative error correction is closely related to the design of a local cellular automaton decoder for the 2D toric code, where each qubit updates its state based only on the states of its neighboring qubits and available syndrome information, following deterministic or probabilistic update rules. This process is inherently parallelizable. Interestingly, a recent work by [10] demonstrates that such a local cellular automaton decoder is indeed feasible, employing a hierarchical construction to address the challenge posed by diffusive quasi-particle excitations. This suggests that Lindblad dynamics incorporating a more complex global jump operator, designed to emulate this cellular automaton decoder, may also enable passive error correction for the 2D toric code.
@@ 1.3 Related works
The thermalization of stabilizer Hamiltonians using a local Davies generator has been explored in several prior works [2, 53, 52, 24, 12, 25]. In [2], the authors demonstrated that the local Davies generator achieves fast thermalization for the 2D toric code. In particular, the spectral gap of the Davies generator, when considering all local Pauli coupling operators, is lower bounded by $\exp(-\Theta(\beta))$. The bound is valid for all temperatures and results in a mixing time (defined via the trace distance) scaling as $t_{\rm mix}=N\exp(\Theta(\beta))$. Along the same direction, [53, 52] considered general stabilizer codes and established a lower bound on the spectral gap using the generalized energy barrier $\overline{\epsilon}$ [52, Definition 13]. Specifically, for any given $\beta>0$, the spectral gap can be lower bounded by $C_{N}\exp\left(-\beta\overline{\epsilon}\right)$, where $C_{N}$ is a technical constant that typically scales as $1/N$. Although these works address thermalization across all temperatures, these lower bounds on the spectral gap are insufficient for efficiently preparing low-temperature thermal states, as the gap decays exponentially in $\beta$. Specifically, when using a local Davies generator to transition between ground states along a local Pauli path, the energy must first increase, requiring the dynamics to overcome the energy barrier to fully thermalize. Furthermore, as analyzed in [52, 37], an exponentially small spectral gap of order $\exp(-\Theta(\beta))$ in a local Davies generator appears inevitable when such energy barriers are present.
In our work, to overcome the bottleneck posed by the energy barrier, we modify the local Davies generator by incorporating appropriate global coupling operators that can directly connect the degenerate ground states, thereby avoiding the issues associated with the generalized energy barrier defined by local Pauli paths. By refining the analysis in [2], we demonstrate that the resulting dynamics exhibit a spectral gap that decays polynomially with the size of the system but remains independent of $\beta$, ensuring fast mixing even at low temperatures. We note that the polynomially decaying spectral gap primarily arises from the mixing rate within the syndrome space, which remains unaffected by the introduction of the new global jump operators acting on the logical space. Consequently, our refined mixing time estimate rigorously demonstrates that the local Davies generator at low temperatures (and even at zero temperature) can efficiently annihilate quasi-particles residing in the syndrome space. A similar phenomenon has been numerically observed in the study of quantum memory [24, 12, 25].